which of the following equations describes the line (-2,4) (-4,-8) check all that apply
a.y-4=2(x+2)
b.y-8=6(x-4)
c.y+4=6(x-2)
d.y+8=6(x+4)
e.y+4=2(x-2)
f.y-4=6(x+2)

Answer by Alan3354(69443)

which of the following equations describes the line (-2,4) (-4,-8) check all that apply
a.y-4=2(x+2)
b.y-8=6(x-4)
c.y+4=6(x-2)
d.y+8=6(x+4)
e.y+4=2(x-2)
f.y-4=6(x+2)
-------------------
Find the slope-point eqn of the line thru the 2 points.
----
Compare to a thru f in the point-slope form
eg
a.y-4=2(x+2)
y = 2x + 8